# Conformal Validity Guarantees Exist for Any Data Distribution

### v2

@@ 5.2 Active Learning Induced MFCS
Now we turn to an active learning setting where MFCS is induced by a max-entropy data acquisition strategy, where the ML agent selects new queries for annotation by maximizing its (Bayesian) predictive uncertainty. In this experiment we use a Gaussian process (GP) regressor as the predictor and take $u_{t}(x)=\mathbb{V}[f(x)|Z_{1:n+t-1}]$ (i.e. the GP posterior variance). See Appendix D.2 for model and dataset details.
@@
Active Learning Experimental Results Figure 3 demonstrates the performance of our proposed MFCS Split Conformal methods for several depths of weight approximation. Whereas the baseline methods lose coverage, the proposed MFCS Split CP methods maintain target coverage for all datasets and all of the 70 evaluated active learning steps. As the active learning process reaches later iterations where more labeled data is available for both training and calibration, the proposed MFCS Split CP methods are still able to maintain target coverage with sharper intervals.
@@ 6 Discussion
In this work we have demonstrated that weighted conformal prediction can theoretically be generalized to any joint data distribution. Exchangeability assumptions are only necessary to rigorously specify the conditions under which a specific practical algorithm (e.g. standard CP) is equivalent to the more general formulation, which is always valid but may not be computable. Promising avenues for future work include extending our general marginal coverage guarantees to more general loss functions (Angelopoulos et al., 2022) and identifying a minimal set of assumptions for good conditional coverage in practice (Gibbs et al., 2023).
We have also provided a practical algorithm for CP in the presence of multiple rounds of agent-induced, data-dependent covariate shift by straightforward derivation, allowing us to maintain coverage on data collected by active learning and black-box optimization agents. Finally we have shown that these agents can attain both empirically valid coverage and finite interval widths by restricting the degree of shift the agent can introduce. These findings bring us to the control/consistency dilemma at the heart of statistical decision-making. To control is to change, but statistics requires consistency. Like the explore/exploit dilemma, the right tradeoff depends on the situation at hand. As machine learning agents assume increasing levels of responsibility in gravely consequential settings, it is more important than ever that we face the control/consistency dilemma head-on.
@@
GitHub repository with code for all experiments: https://github.com/drewprinster/conformal-mfcs
@@ Appendix A Main Result Proof and Details
In this section we present the proof for our main result given in Theorem (4.1). Our proof builds on ideas in Tibshirani et al. (2019)’s alternate proof of their Lemma 1 (for conformal prediction assuming exchangeable data) and their proof of their Lemma 3 (for conformal prediction assuming weighted exchangeable data). However, those proofs make key substitutions using the definitions of exchangeability and weighted exchangeability respectively, which thereby yields guarantees that are premised on those definitions. In contrast, the proof we present here makes no such substitution, and thus holds for an arbitrary joint probability density function $f$. We will call attention to the key step where our proof critically differs from those in Tibshirani et al. (2019) in substance, and even for steps that we have in common with Tibshirani et al. (2019) we aim to provide more explicit derivation explanations for increased accessibility (e.g., where explicit step justifications may be omitted in Tibshirani et al. (2019)).
As we sketch in Section 4.1 of our main text, our proof corresponds to a view of CP as an inverted permutation test. Let $E_{z}$ denote the event of observing the unordered set [^9] As in Tibshirani et al. (2019) we assume for simplicity that the score values $V_{1},...,V_{n+1}$ are distinct almost surely, though the result also holds in more general cases (but with more complicated notation involving randomness from uniform random variables). of our data values. In particular, $E_{z}$ denotes the event that $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$, but importantly, the lack of ordering means that this does not imply that $Z_{i}=z_{i}$ for all $i\in\{1,...,n+1\}$; we do not yet know whether $z_{1}$ is the value obtained by $Z_{1}$, or by $Z_{2}$, or by $Z_{n+1}$, and so on. Similarly, let $E_{v}$ denote the event $\{V_{1},...,V_{n+1}\}=\{v_{1},...,v_{n+1}\}$ for the scores. Also, recall the simplifying notation $v_{i}=\mathcal{\widehat{S}}(z_{i})$ for the value computed by our fitted score function $\mathcal{\widehat{S}}$ with input $z_{i}$ (where $\mathcal{\widehat{S}}(z_{i})=\mathcal{S}(z_{i},z_{1:(n+1)})$, that is $\mathcal{\widehat{S}}$ is fit on the set $z_{1:(n+1)}=\{z_{1},...,z_{n+1}\}$, the second input to $\mathcal{S}$).
@@
which follows since, conditioned on $E_{z}$, there are an equal number of scores $v_{1},...,v_{n+1}$ as data points $z_{1},...,z_{n+1}$, thus implying that the fitted (symmetric) score function $\mathcal{\widehat{S}}$ induces a bijection between the discrete set of distinct data observations and the discrete set of distinct scores. [^10] To be thorough and precise, the equation $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ relies on the convention in conformal prediction that $\mathcal{S}$ is a symmetric score function (see Remark 4.3). We elaborate on this point in Appendix A.3. Then, using the definition of conditional probability and the law of total probability (LOTP), it follows that
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ =$=\frac{\mathbb{P}\{Z_{n+1}=z_{i},E_{z}\}}{\mathbb{P}\{E_{z}\}}$
@@
where the last step follows by applying the LOTP in both the numerator and the denominator. For example, in the denominator the summation is over all possible permutations $\sigma$ of the value, i.e., all the different orderings in which the event $E_{z}$ could have occurred. Even more explicit notation could be to write out $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})$, to emphasize that the permutation is on the values and not on the random variables in the joint distribution $f$.
Here is where our proof differs critically from the analogous proofs in Tibshirani et al. (2019). At this point, their proof corresponding to exchangeable conformal prediction uses exchangeability to substitute in $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$, since under exchangeability $f$ is invariant to permutations $\sigma$; meanwhile, their proof corresponding to weighted conformal prediction substitutes in $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1},...,z_{n+1})$, since under their definition of weighted exchangeability $f$ factorizes into weight functions $w_{j}$ and a core function $g$ that is invariant to $\sigma$. Here, unlike in Tibshirani et al. (2019), we instead maintain the generality of Eq. (10) to prove our more general conformal prediction result that holds for any $f$.
@@
and marginalizing yields
@@
We note that statement (11) can be viewed as a generalized version of both Lemma 1 and Lemma 3 in Tibshirani et al. (2019). We now connect statement (11) to the general conformal prediction set defined in Theorem 4.1 to obtain the result. To do so, we observe that
@@
where the right-hand side is obtained from the left-hand side by replacing $\delta_{V_{n+1}}$ with $\delta_{\infty}$ to conservatively increase the value of the right side relative to the left. Now, with the notation $V_{i}=V_{i}^{(X_{n+1},Y_{n+1})}$ for $i\in\{1,...,n+1\}$, note that the right hand side of (12) is the same quantile as is used in the construction of the general conformal prediction interval in 4.1. That is, $Y_{n+1}\in\widehat{C}_{n}(X_{n+1})$ if and only if
@@
Due to statement (12), the probability of the event (13) is thus lower bounded by statement (11). Letting $\beta=1-\alpha$ in the quantile function then gives the result.
@@
In Theorem 4.1 and the proof in Appendix A.1 we use score functions that correspond to the “ordinary” full conformal method Vovk et al. (2005). However, the same result and proof technique would also apply to generalizing the “deleted” or “leave-one-out” full conformal method (which is computationally more expensive than ordinary full CP but tends to obtain more informative prediction sets by avoiding overfitting in its construction) (Vovk et al., 2005) or the split conformal method (Papadopoulos, 2008). The score functions for the deleted full CP method would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{-i}\cup\{(x,y)\})$ for $i\in\{1,...,n\}$, where $Z_{-i}=Z_{1:n}\backslash Z_{i}$, and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:n}\big{)}$. For split CP, the score functions could be defined using independent training and calibration sets. Letting $Z^{0}_{1:m}=\{Z_{i}^{0}\}_{i=1}^{m}$ denote the training data and $Z_{1:n}=\{Z_{i}\}_{i=1}^{n}$ denote the calibration data, the split CP score functions would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{1:m}^{0})$ for $i\in\{1,...,n\}$ and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:m}^{0}\big{)}$.
@@ Appendix B Deriving Weighted Conformal Validity Guarantees and Algorithms for Data under MFCS
We begin by restating the formal assumptions we used to characterize MFCS:
@@
Recall that informally, these assumptions are stating that (after an IID initialization of $n$ datapoints) the distribution of the covariates $X$ at time $t$ can change arbitrarily depending on past observations $Z_{1:(n+t-1)}=\{Z_{1},...,Z_{n+t-1}\}$, while the conditional label distribution $Y\mid X$ is assumed to remain invariant.
@@
$=\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t}=z_{n+t})$ More explicit notation
$=\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1},X_{n+t}=x_{n+t},Y_{n+t}=y_{n+t})$ $Z_{n+t}=(X_{n+t},Y_{n+t})$
$\ \begin{matrix}[l]=\mathbb{P}(Y_{n+t}=y_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1},X_{n+t}=x_{n+t})\\ \qquad\cdot\ \mathbb{P}(X_{n+t}=x_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\\ \qquad\cdot\ \mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\end{matrix}$ Chain rule or conditional probability definition
$\ \begin{matrix}[l]=\mathbb{P}(Y_{n+t}=y_{n+t}\ |\ X_{n+t}=x_{n+t})\\ \qquad\cdot\ \mathbb{P}(X_{n+t}=x_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\\ \qquad\cdot\ \mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\end{matrix}$ $\begin{matrix}\text{MFCS invariance assumption:}\\ Y\ |\ X\text{ invariant }\implies\\ (Y_{n+t}\ |\ X_{n+t})\perp\!\!\!\perp Z_{1},...,Z_{n+t-1}\end{matrix}$
These steps are not specific to $n+t$, however; we can repeat the same procedure for $\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})$ and the index $n+t-1$, and so on. When these steps are performed for all indices $i\in\{1,...,n+t\}$, we obtain
$f$ $(z_{1},...,z_{n+t})$
$=\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(X_{j}=x_{j}\ |\ Z_{1}=z_{1},...,Z_{j-1}=z_{j-1})\cdot\mathbb{P}(Y_{j}=y_{j}|X_{j}=x_{j})\Big{]}$
$=\prod_{j=1}^{n+t}\Big{[}\underbrace{\mathbb{P}(X_{j}=x_{j}\ |\ Z_{1}=z_{1},...,Z_{j-1}=z_{j-1})}_{\begin{matrix}\text{Time-dependent factors}\end{matrix}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(Y_{j}=y_{j}|X_{j}=x_{j})\Big{]}}_{\begin{matrix}\text{Time-invariant factor}\end{matrix}},$
that is, we obtain the factorization into time-dependent and time-invariant factors provided in the main paper Eq. (7).
@@
$\mathbb{P}$$\{Z_{n+t}=z_{i}\mid E_{z}^{(t)}\}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}f(z_{\sigma(1)},...,z_{\sigma(n+t)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+t)})}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(Y_{j}=y_{\sigma(j)}\ |\ X_{j}=x_{\sigma(j)})\Big{]}}{\sum_{\sigma}\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(Y_{j}=y_{\sigma(j)}\ |\ X_{j}=x_{\sigma(j)})\Big{]}}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})},$
where this result is equivalent to Eq. (4.3) in the main paper; it follows because $\prod_{j=1}^{n+t}\mathbb{P}(Y_{j}=y_{\sigma(j)}\ |\ X_{j}=x_{\sigma(j)})$ is invariant to permutations $\sigma$, and thus cancels out in the ratio. For more concise notation to refer to these weights derived using MFCS assumptions, let us define
$\mathbb{P}_{n+t}^{(\text{MFCS})}\{z_{i}\mid E_{z}^{(t)}\}:=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}.$ (14)

### v3

@@ 5.2 Active Learning-Induced MFCS
Now we turn to an active learning setting where MFCS is induced by a max-entropy data acquisition strategy, where the ML agent selects new queries for annotation by maximizing its (Bayesian) model uncertainty. In this experiment we use a Gaussian process (GP) regressor as the predictor and take $u_{t}(x)=\mathbb{V}(\mu(x)\mid Z_{\text{train}}^{(t)})$ (i.e. the GP posterior variance). See Appendix E.2 for model and dataset details.
@@
Active Learning Experimental Results Figure 3 demonstrates the performance of our proposed MFCS Split Conformal methods for several depths of weight approximation. Whereas the baseline methods lose coverage, the proposed MFCS Split CP methods maintain target coverage for all datasets and across the 70 evaluated active learning steps. As the active learning process reaches later iterations where more labeled data is available for both training and calibration, the proposed MFCS Split CP methods are still able to maintain target coverage with sharper intervals.
@@ 6 Discussion
In this paper we have demonstrated how weighted conformal prediction can theoretically generalize to any joint data distribution. Exchangeability conditions are only necessary to rigorously specify when a specific practical algorithm (e.g., standard CP) is equivalent to the more general formulation, which is always valid but may not be computable. Promising future directions include further addressing practical bottlenecks, extending our general marginal coverage guarantees to other loss functions (Angelopoulos et al., 2022), and identifying minimal assumptions for strong conditional coverage in practice (Barber et al., 2021; Gibbs et al., 2023).
We have also provided practical CP algorithms for multiple rounds of agent-induced, data-dependent covariate shifts, which maintains coverage on data collected by active learning and black-box optimization agents. Finally, we have shown that these agents can attain both empirically valid coverage and informative intervals by restricting the degree of shift introduced (Appendix C). These findings bring us to the control-consistency dilemma at the heart of statistical decision-making. To control is to change, but statistics requires consistency. Like the explore-exploit dilemma, the right tradeoff depends on the situation at hand. As machine learning agents assume increasing levels of responsibility in gravely consequential settings, it is more important than ever that we face the control-consistency dilemma head-on.
@@
Repository with code for all experiments: https://github.com/drewprinster/conformal-mfcs
@@ Appendix A Main Result Proof and Details
In this section we present the proof for our main result given in Theorem (4.1). Our proof builds on ideas in Tibshirani et al. (2019)’s alternate proof of their Lemma 1 (for conformal prediction assuming exchangeable data) and their proof of their Lemma 3 (for conformal prediction assuming weighted exchangeable data). However, those proofs make key substitutions using the definitions of exchangeability and weighted exchangeability respectively, which thereby yields guarantees that are premised on those definitions. In contrast, the proof we present here makes no such substitution, and thus holds for an arbitrary joint probability density function (PDF) [^9] More generally, $f$ can be a valid Radon-Nikodym derivative with respect to an arbitrary base measure: in the discrete case $f$ is a probability mass function (PMF); it could also be a mixture of PDFs and PMFs. $f$. We will call attention to the key step where our proof critically differs from those in Tibshirani et al. (2019) in substance, and even for steps that we have in common with Tibshirani et al. (2019) we aim to provide more explicit derivation explanations for increased accessibility (e.g., where explicit step justifications may be omitted in Tibshirani et al. (2019)).
As we sketch in Section 4.1 of our main text, our proof corresponds to a view of CP as an inverted permutation test. Let $E_{z}$ denote the event of observing the unordered set [^10] As in Tibshirani et al. (2019) we assume for simplicity that the score values $V_{1},...,V_{n+1}$ are distinct almost surely, though the result also holds in more general cases (but with more complicated notation involving randomness from uniform random variables to break ties). of our data values. In particular, $E_{z}$ denotes the event that $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$, but importantly, the lack of ordering means that this does not imply that $Z_{i}=z_{i}$ for all $i\in\{1,...,n+1\}$; we do not yet know whether $z_{1}$ is the value obtained by $Z_{1}$, or by $Z_{2}$, or by $Z_{n+1}$, and so on. Similarly, let $E_{v}$ denote the event $\{V_{1},...,V_{n+1}\}=\{v_{1},...,v_{n+1}\}$ for the scores. Also, recall the simplifying notation $v_{i}=\mathcal{\widehat{S}}(z_{i})$ for the value computed by our fitted score function $\mathcal{\widehat{S}}$ with input $z_{i}$ (where $\mathcal{\widehat{S}}(z_{i})=\mathcal{S}(z_{i},z_{1:(n+1)})$, that is $\mathcal{\widehat{S}}$ is fit on the set $z_{1:(n+1)}=\{z_{1},...,z_{n+1}\}$, the second input to $\mathcal{S}$).
@@
which follows since, conditioned on $E_{z}$, there are an equal number of scores $v_{1},...,v_{n+1}$ as data points $z_{1},...,z_{n+1}$, thus implying that the fitted (symmetric) score function $\mathcal{\widehat{S}}$ induces a bijection between the discrete set of distinct data observations and the discrete set of distinct scores. [^11] To be thorough and precise, the equation $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}=\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ relies on the convention in conformal prediction that $\mathcal{S}$ is a symmetric score function (see Remark 4.3). We elaborate on this point in Appendix A.3. Then, using the definition of conditional probability and the law of total probability (LOTP), it follows that
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ $=\frac{p\{Z_{n+1}=z_{i},E_{z}\}}{p\{E_{z}\}}$
@@
where the last step follows by applying the LOTP in both the numerator and the denominator. For example, in the denominator the summation is over all possible permutations $\sigma$ of the value, i.e., all the different orderings in which the event $E_{z}$ could have occurred. Even more explicit notation could be to write out $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f_{Z_{1},...,Z_{n+1}}(z_{\sigma(1)},...,z_{\sigma(n+1)})$ if $f$ is a PDF, or $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})$ if $f$ is a probability mass function (PMF), to emphasize that the permutation is on the observed values and not on the random variables in the joint distribution.
Here is where our proof differs critically from the analogous proofs in Tibshirani et al. (2019). At this point, their proof corresponding to exchangeable conformal prediction uses exchangeability to substitute in $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$, since under exchangeability $f$ is invariant to permutations $\sigma$ (see Appendix D.1 for details); meanwhile, their proof corresponding to weighted conformal prediction substitutes in $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1},...,z_{n+1})$, since under their definition of weighted exchangeability $f$ factorizes into weight functions $w_{j}$ and a core function $g$ that is invariant to $\sigma$ (see Appendix D.2 for details on weighted exchangeability). Here, unlike in Tibshirani et al. (2019), we instead maintain the generality of Eq. (10) by avoiding any substitutions, and this will allow us to prove the more general conformal prediction result, which holds for any $f$.
@@
and marginalizing over draws of the event $E_{z}$ yields
@@
We note that statement (11) can be viewed as a general quantile lemma, which generalizes both Lemma 1 and Lemma 3 in Tibshirani et al. (2019). We now connect statement (11) to the general conformal prediction set defined in Theorem 4.1 to obtain the result. To do so, we observe that
@@
where the right-hand side is obtained from the left-hand side by replacing $\delta_{V_{n+1}}$ with $\delta_{\infty}$ to conservatively increase the value of the right side relative to the left. Now, with the notation $V_{i}=V_{i}^{(X_{n+1},Y_{n+1})}$ for $i\in\{1,...,n+1\}$, note that the right hand side of (12) is the same quantile as is used in the construction of the general conformal prediction interval in 4.1. That is, $Y_{n+1}\in\widehat{\mathcal{C}}_{n}(X_{n+1})$ if and only if
@@
Due to statement (12), the probability of the event (13) is thus lower bounded by statement (11). Setting $\beta=1-\alpha$ in the quantile function then gives the result.
@@
In Theorem 4.1 and the proof in Appendix A.1 we use score functions that correspond to the “ordinary” full conformal method Vovk et al. (2005). However, the same result and proof technique would also apply to generalizing the “deleted” or “leave-one-out” full conformal method (which is computationally more expensive than ordinary full CP but can obtain more informative prediction sets by avoiding overfitting in its construction) (Vovk et al., 2005) or the split conformal method (Papadopoulos, 2008). The score functions for the deleted full CP method would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{-i}\cup\{(x,y)\})$ for $i\in\{1,...,n\}$, where $Z_{-i}=Z_{1:n}\backslash Z_{i}$, and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:n}\big{)}$. For split CP, the score functions could be defined using independent training and calibration sets. Letting $Z^{0}_{1:m}=\{Z_{i}^{0}\}_{i=1}^{m}$ denote the training data and $Z_{1:n}=\{Z_{i}\}_{i=1}^{n}$ denote the calibration data, the split CP score functions would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{1:m}^{0})$ for $i\in\{1,...,n\}$ and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:m}^{0}\big{)}$.
@@ Appendix B Deriving Weighted Conformal Validity Guarantees and Algorithms for Data under MFCS
Regarding notation: We will use $f$ to denote the joint probability density function (PDF) for the sequence of random variables $Z_{1},...,Z_{n+t}$ and $p$ to denote more specific (conditional) PDFs, for example $p_{Y|X}=dP_{Y|X}$. To lighten notation we will sometimes omit implicit random variables; for example, we will write $p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\mid Z_{1}=z_{1},...,Z_{j-1}=z_{j-1})$ equivalently as $p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\mid z_{1},...,z_{j-1})$ or even as $p(x_{j}\mid z_{1},...,z_{j-1})$, if $X_{j}|Z_{1},...,Z_{j-1}$ are clear from context and if the shorthand may help declutter the notation. [^13] We also use $p(x\mid Z_{1},...,Z_{j-1})$ to denote the generic the conditional density function that can change depending on different realized values of $Z_{1},...,Z_{j-1}$ (e.g., it could be that $p(x\mid Z_{1}=z_{1},...,Z_{j-1}=z_{j-1})\neq p(x\mid Z_{1}=z_{1}^{\prime},...,Z_{j-1}=z_{j-1}^{\prime})$ for some $(z_{1},...,z_{j-1})\neq(z_{1}^{\prime},...,z_{j-1}^{\prime})$), whereas $p(x\mid z_{1},...,z_{j-1})$ denotes the specific conditional density, conditional on the joint event $Z_{1}=z_{1},...,Z_{j-1}=z_{j-1}$.
@@
We begin by restating the formal assumptions we used to characterize MFCS in Eq. (3):
@@
Recall that informally, these assumptions are stating that (after an IID initialization of $n$ datapoints) the distribution of the covariates $X$ at time $t$ can change arbitrarily depending on past observations $Z_{1:(n+t-1)}=\{Z_{1},...,Z_{n+t-1}\}$, while the conditional label distribution $Y\mid X$ is assumed to remain invariant. Note that $Y\mid X$ being invariant implies $Z_{1},...,Z_{n+t-1}\perp\!\!\!\perp Y_{n+t}\mid X_{n+t}$, where the latter can be read as “$Y_{n+t}$ is independent of past observations $Z_{1},...,Z_{n+t-1}$, given $X_{n+t}$,” and moreover that $Y_{i}|X_{i}\stackrel{{\scriptstyle\mathclap{d}}}{{=}}Y|X$ for all $i\in\{1,...,n+T\}$.
@@
$=f(z_{1},...,z_{n+t-1},x_{n+t},y_{n+t})$ $z_{n+t}=(x_{n+t},y_{n+t})$
$\ \begin{matrix}[l]=\ p_{Z_{1},...,Z_{n+t-1}}(z_{1},...,z_{n+t-1})\cdot p_{X_{n+t}|Z_{1},...,Z_{n+t-1}}(x_{n+t}\mid z_{1},...,z_{n+t-1})\\ \qquad\cdot\ p_{Y_{n+t}|Z_{1},...,Z_{n+t-1},X_{n+t}}(y_{n+t}\mid z_{1},...,z_{n+t-1},x_{n+t})\end{matrix}$ $\begin{matrix}\text{Chain rule of probability or}\\ \text{conditional probability density def.}\end{matrix}$
$\ \begin{matrix}[l]=\ p_{Z_{1},...,Z_{n+t-1}}(z_{1},...,z_{n+t-1})\cdot p_{X_{n+t}|Z_{1},...,Z_{n+t-1}}(x_{n+t}\mid z_{1},...,z_{n+t-1})\\ \qquad\cdot\ p_{Y|X}(y_{n+t}\mid x_{n+t})\end{matrix}$ $\begin{matrix}\text{MFCS invariance assumption:}\\ Y\ |\ X\text{ invariant }\implies\\ Z_{1},...,Z_{n+t-1}\perp\!\!\!\perp Y_{n+t}\mid X_{n+t}\\ \text{and }Y_{n+t}|X_{n+t}\stackrel{{\scriptstyle\mathclap{d}}}{{=}}Y|X\end{matrix}$
These steps are not specific to $n+t$, however; we can repeat the same procedure for $p_{Z_{1},...,Z_{n+t-1}}(z_{1},...,z_{n+t-1})$ and the index $n+t-1$, and so on. When these steps are performed for all indices $i\in\{1,...,n+t\}$, we obtain
$f(z_{1},...,z_{n+t})$ $=\prod_{j=1}^{n+t}\Big{[}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\mid z_{1},...,z_{j-1})\cdot p_{Y|X}(y_{j}|x_{j})\Big{]}$
$=\prod_{j=1}^{n+t}\Big{[}\underbrace{p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{j}\mid z_{1},...,z_{j-1})}_{\begin{matrix}\text{Time-dependent factors}\end{matrix}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{j}|x_{j})\Big{]}}_{\begin{matrix}\text{Time-invariant factor}\end{matrix}},$
that is, we obtain the factorization into time-dependent and time-invariant factors provided in the main paper Eq. (7) (where in the main paper we omitted the random variables from the subscripts for ligher notation).
@@
$\mathbb{P}\{Z_{n+t}=z_{i}\mid E_{z}^{(t)}\}$ $=\frac{\sum_{\sigma:\sigma(n+t)=i}f(z_{\sigma(1)},...,z_{\sigma(n+t)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+t)})}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\Big{[}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{\sigma(j)}\mid x_{\sigma(j)})\Big{]}}{\sum_{\sigma}\prod_{j=1}^{n+t}\Big{[}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}p_{Y|X}(y_{\sigma(j)}\mid x_{\sigma(j)})\Big{]}}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})},$
where this result is equivalent to Eq. (4.3) in the main paper; it follows because $\prod_{j=1}^{n+t}p_{Y|X}(y_{\sigma(j)}\mid x_{\sigma(j)})$ is invariant to permutations $\sigma$, and thus cancels out in the ratio. For more concise notation to refer to these weights derived using MFCS assumptions, let us define
$\mathbb{P}_{n+t}^{(\text{MFCS})}\{z_{i}\mid E_{z}^{(t)}\}:=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}p_{X_{j}|Z_{1},...,Z_{j-1}}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}.$ (14)